If one-half a number is subtracted from five-sixths of the number, the difference is 6. Find the number.
Answer by ankor@dixie-net.com(22740)

:
Let x = the number:
:
If one-half a number is subtracted from five-sixths of the number, the difference is 6.
(5/6)x - (1/2)x = 6
:
Find the number.
Get rid of the denominators, multiply equation by 6, resulting in:
5x - 3x = 36
2x = 36
x = 36/2
x = 18
:
:
Check solution x = 18:
(5/6)18 - (1/2)18 =
15 - 9 = 6